### v1

# EMF Mitigation via 5G and 6G Scheduling

Abstract
High antenna directivity allows for high throughput transmission but also increases the exposure to electromagnetic field (EMF) of the end-users. Health regulations impose limitations on the incident power density, that generate a negative impact on network performance.
In this work we focus at the slot-by-slot operations of a cellular Medium Access Control (MAC) scheduler to constrain the short-term EMF exposure upon real-time resource allocation, minimizing the impacts on network performance. We assume that the long-term EMF exposure is controlled by a proper outer-loop technique, that is not the object of this paper. Due to the minimal computational complexity allowed in MAC scheduling, existing solutions allowing practical implementation are few and focused at sub-optimal approaches curbing radio resource allocation.
Our contribution is the derivation of a computationally efficient water-filling solution to allocate power and - then - resources, with a feasible integration of the necessary algorithms in the operations of a 5G MAC scheduler. We finally evaluate our proposal versus the prior art approaches with system level simulations with realistic modeling of physical and MAC level cellular procedures. We conclude that our proposal can control EMF with considerable less impact on network performance, making it a standout candidate for 5G and future 6G MAC scheduler implementations.
Index Terms:
EMF, RF exposure, EIRP, MAC Scheduling, Power Control.
I Introduction
The high beam directivity characterizing the deployment of fifth generation (5G) cellular networks has sparked concerns regarding potential health risks associated with electromagnetic field (EMF) exposure. However, regulators and the research community have long been committed to thoroughly understanding and addressing any health implications of EMF exposure through continuous regulation and oversight efforts.
The guidelines by the International Commission on Non-Ionizing Radiation Protection (ICNIRP) [1] specify the EMF exposure limits, depending on the operating radio frequency and enforced on the average power density over a sliding time window ranging from 1 to 30 minutes [1, 2]. Such limits can also be defined country-wise.
Ensuring the effective enforcement of such limits while minimizing the impact on the network performance will become increasingly challenging in sixth generation (6G) Multiple Inputs Multiple Outputs (MIMO) deployments where hundreds or thousands antenna elements, hence even higher beam gains than in 5G, are expected [3].
By adopting a conservative approach, one can ensure that EMF constraints are met even in the worst case scenario where users consistently find themselves in the direction of maximum beam gain. This translates to limiting the Effective Isotropic Radiated Power (EIRP) of the radiating antenna array.
Arguably, imposing real-time constraints on EIRP in cellular networks is no easy task, due to the scheduler complexity and the tight computation latency constraint. Existing approaches [4, 5] propose power control procedures limiting the usable bandwidth; [6] optimizes precoding for MIMO systems under the assumption of full channel knowledge. However, this assumption is unfeasible in reality as it would require heavy inter-working between physical and Medium Access Control (MAC) layer. Finally, it has been shown that imposing EIRP constraints affects network performance [7], highlighting the importance of designing algorithms that minimize the performance degradation while being implementable in MAC layer real-time procedures.
Our contribution. We consider the framework illustrated in Fig. 1, where an “outer-loop” mechanism operates on a slow time scale, monitoring the average EIRP over a sliding window of a few minutes, as prescribed by ICNIRP. Such mechanism sets the EIRP budget for the next “period” which encompasses the next few hundred of slots. The design of the outer-loop mechanism is not the focus of this paper and can be found, e.g., in [8]. In this paper we focus on the “inner-loop” aspect depicted in Fig. 1, which enforces the EIRP budget in real-time via MAC scheduling operations. Initially, the subset of User Equipments (UEs) eligible for transmission in the current slot is selected as described, e.g., in [9]. Then, the EIRP is controlled by allocating Physical Resource Blocks (PRBs), power and Modulation and Coding Scheme (MCS) fairly across UEs. Our proposal comprises two subroutines: the first spreads the EIRP budget across slots within the same period, to accommodate traffic bursts while avoiding budget depletion before the period ends. The second subroutine performs a fair allocation across the UEs while fulfilling the constraint on the per-slot EIRP budget.
Prior art [4, 5] proposed constraining resources solely in the frequency - hence PRB- domain to limit EIRP. Instead, we propose dynamically adjusting the user MCS and power to maximize the global UE throughput fairness while adhering to an EIRP constraint. The concept of reducing MCS and power while spreading the transmission across more radio time and frequency resources was already introduced in [10, 11]. Yet, constraints on EIRP were not considered since the primary focus was mitigating interference in non-full buffer scenarios.
Our intuition suggests that, to control EIRP, reducing power similarly to the approach in [10, 11] should be preferred over reducing bandwidth. This reasoning stems directly from the classic Shannon formula: reducing bandwidth leads to a linear decrease in both rate and EIRP, while decreasing power results in a logarithmic reduction in rate. These effects tend to coincide in the low SINR regime. Therefore, we expect power-based EIRP control solutions to drastically outperform bandwidth based approaches, if properly implemented. As a side-product, overall network interference will be reduced, as demonstrated in [10, 11].
We evaluate our proposal in a highly realistic 3rd Generation Partnership Project (3GPP) calibrated simulation setup, with a full physical and MAC implementation of a MIMO cellular scenario. The results enable us to draw conclusions regarding the potential architecture and algorithms for deployment in a MAC scheduler of a g-Node B (gNB) in cellular networks to control EIRP with minimal impact on system performance.
[…]
Figure 1: Block diagram of EIRP control operations considered in this work. Slot index $k$ has been omitted in inner-loop operations for readability. The main focus of our work concerns the red blocks, while the other parts are either legacy MAC scheduling operations in the inner-loop, or outer-loop operations discussed in [8]
II EMF exposure model
We address the downlink (DL) scheduling problem of allocating PRBs and power to users while mitigating the human exposure to EMF according to the health regulations imposed by the ICNIRP.
We first introduce some notation. We let $A_{u,k}$ and $P_{u,k}$ be the number of allocated PRBs and the transmit power per PRB for user $u$ in slot $k$, respectively. We denote by $\beta_{u,k}$ the beam serving user $u$ in slot $k$ and we call $G_{u,k}(\phi,\theta)$ the antenna gain in azimuth/elevation direction $(\phi,\theta)$ when beam $\beta_{u,k}$ is used. The set of active UEs, i.e., the UEs with DL data in the buffer, is $\mathcal{U}_{k}$.
The EIRP emitted by a radiating antenna array in the azimuth/elevation direction $(\phi,\theta)$ in slot $k$ writes:
$\mathrm{EIRP}_{k}(\phi,\theta):=\sum_{u\in\mathcal{U}_{k}}A_{u,k}\,P_{u,k}\,G_{u,k}(\phi,\theta).$ (1)
The power density $D$ measured at distance $R$ from the antenna in the direction $(\phi,\theta)$ is proportional to the EIRP, and equals $D=\frac{\mathrm{EIRP}(\phi,\theta)}{4\pi R^{2}}$. ICNIRP imposes health-based limits on $D$ [1], which can be translated into an upper bound $\overline{C}$ on the average EIRP, depending on the statistical distribution of the distance of humans from the antenna [7]. Depending on the presence and type of human activities present in the cell (schools, hospitals, etc.), different limits may need to be imposed in different directions. To account for this, we partition the set of all azimuth/elevation angles, defined with respect to the radiating antenna, into a set of segments segments $\mathcal{S}$, and assign a specific EIRP limit $\overline{C}_{s}$ to each segment $s\in\mathcal{S}$. We then define the EIRP consumption $c_{k,s}$ as the maximum EIRP in segment $s$:
$c_{k,s}:=$ $\,\max_{(\phi,\theta)\in s}\mathrm{EIRP}_{k}(\phi,\theta).$ (2)
We call period $t$ a set of $K\geq 1$ consecutive slots starting from $k_{0}^{t}$, and we define $c^{t}_{s}:=\sum_{k=k_{0}^{t}}^{k_{0}^{t}+K-1}c_{k,s}$ the sum of EIRP consumption over period $t$. To fulfill the ICNIRP health regulations we impose that the average EIRP consumption over every sliding window of $W$ consecutive periods does not exceed $\overline{C}_{s}$, for each segment $s$:
$\frac{1}{W}\sum_{i=0}^{W-1}c^{t-i}_{s}\leq\overline{C}_{s},\qquad\forall\,t,\ s\in\mathcal{S}.$ (3)
III Power control for EMF exposure mitigation
As illustrated in Fig. 1, an outer-loop algorithm computes the limit $\gamma_{s}^{t}$ on the EIRP consumption in period $t$, such that
$0\leq c^{t}_{s}\leq\gamma^{t}_{s},\qquad\forall\,t,\ s\in\mathcal{S}.$ (4)
The design of the outer-loop algorithm is not the focus of this paper and is discussed, e.g., in [8]. Then, we consider $\gamma^{t}_{s}$ as a predefined input for each $s,t$. In this paper we study how to enforce the EIRP cap (4) on a per-slot and on a per-user basis via a so-called inner-loop control mechanism within DL MAC scheduler operations depicted in Fig. 1.
Without loss of generality we consider $k_{0}^{t}=1$. We then split our problem into two sub-problems that we solve separately.
  1. 1.
    Per-slot EIRP constraint. For each segment $s$ and in each period $t$, it must hold that $\sum_{k=1}^{K}c_{k,s}\leq\gamma^{t}_{s}$. We smooth out the EIRP consumption over time by computing an upper limit - or slot budget - $b_{k,s}$ for $c_{k,s}$, for each slot $k$ and segment $s$.
  2. 2.
    Power allocation. At slot $k$, we allocate power fairly across UEs by ensuring $c_{k,s}\leq b_{k,s}$.
For simplicity of notation, we herafter drop the dependence of variables on period $t$.
III-A Per-slot EIRP constraint design
Let us fix the segment $s$. At the beginning of each slot $k=1,\dots,K$, given the past consumption $c_{1,s},\dots,c_{k-1,s}$, we must cap the EIRP consumption for the upcoming slot to a value $b_{k,s}$. We propose to inject a portion $(1-\epsilon_{s})$ of the period’s budget $\gamma_{s}$ at slot $k=0$ and spread the remainder by injecting $\epsilon_{s}\gamma_{s}/K$ at each slot, with $0\leq\epsilon_{s}\leq 1$. Therefore, the available slot budget at slot $k$ is
$b_{k,s}=\max\left(\left(1-\epsilon_{s}\frac{K-k}{K}\right)\gamma_{s}-\sum_{i=1}^{k}c_{i,k},0\right).$ (5)
The difficulty in designing the portion $\epsilon_{s}$ resides in the fact that future user requests are unknown. On the one hand, we want to avoid to greedily deplete the budget early on in the period with $\epsilon_{s}$ close to zero, which would lead to empty slots later on. On the other hand, we want to refrain from introducing unnecessarily delay by curbing consumption in an overly conservative manner with $\epsilon_{s}$ close to one.
Similarly, we would like to avoid the possible excessively low slot budget as computed in (5). Therefore, the slot budget can be lower bounded by $\rho_{s}^{*}c_{s}^{*}$, where $c_{s}^{*}$ is the maximum EIRP that can be radiated in the segment $s$ in one slot on all available PRBs and full power and maximum beam gain, and $0<\rho_{s}^{*}c_{s}^{*}<\gamma_{s}/K$.
Optionally, we propose to start curbing EIRP once the slot budget is below a certain guard threshold $b_{s}^{*}$, controlling EIRP emission before the slot budget gets too low. Therefore, we also study the following slot budget refinement
$b^{\prime}_{k,s}=c_{s}^{*}\exp\left(\ln(\rho_{s}^{*})\max({1-{b_{k,s}}/{b_{s}^{*}}},0)\right).$ (6)
III-B Power Limiting (PL) techniques
We now show how to enforce the EIRP constraint $b_{k,s}$ on segment $s$ by allocating power to active UEs in a fair manner. We assume that $A_{u,k}$ PRBs have been pre-allocated to each active UE $u\in\mathcal{U}_{k}$ as a function of—amongst others—the amount of bits present in the buffer for each UE, the selected MCS and the corresponding rate $m_{u},r_{u}$, respectively. In our preferred implementation, this is performed via the method described in [11], which downgrades the UEsMCS and reduces the power accordingly, hence increasing the required PRB allocation, but without exceeding the PRBs available in the current slot. As a result, the transmit power and resulting interference in neighbouring cells are reduced.
Two simplifying approximations. Since power allocation must be decided at every slot, computational complexity is a bottleneck. To simplify the problem we first approximate that each beam $\beta$ has positive gain only in one segment $s(\beta)$, i.e., the one containing its main lobe. We call $\mathcal{U}_{k,s}\subset\mathcal{U}_{k}$ the set of active users served by a beam whose main lobe is in segment $s$, i.e., $\mathcal{U}_{k,s}:=\{u\in\mathcal{U}_{k}:\,s(\beta_{u,k})=s\}$. Then, the EIRP consumption $c_{k,s}$ depends only on users $\mathcal{U}_{k,s}$:
$c_{k,s}=\max_{(\phi,\theta)\in s}\sum_{u\in\mathcal{U}_{k,s}}A_{u,k}\,P_{u,k}\,G_{u,k}(\phi,\theta)$ (7)
and the power allocation $P_{u,k}$ for a user in $\mathcal{U}_{k,s}$ does not impact the EIRP consumption over a different segment $s^{\prime}\neq s$. In this case, we can optimize the user power allocation independently for each segment.
Next, computing $c_{k,s}$ as in (7) requires the solution of an optimization problem. To circumvent it and simplify computations, we pre-compute the maximum gain for each beam:
$\widehat{G}_{u,k}:=\max_{(\phi,\theta)\in s}G_{u,k}(\phi,\theta)$ (8)
and we bound $c_{k,s}$ via a simple-to-compute expression:
$c_{k,s}\leq\,\sum_{u\in\mathcal{U}_{k,s}}A_{u,k}\,P_{u,k}\,\widehat{G}_{u,k}.$ (9)
Note that equality in (9) holds only if the gain is maximal in the same direction for all beams serving users in $\mathcal{U}_{k,s}$.
Optimization problem. We now formulate the problem of power allocation across UEs. For notation simplicity we drop the variable dependence on $s,k$, with the understanding that the same procedure is used at each slot $k$ and segment $s$.
Let $r_{u}(P_{u})$ be the function mapping the power allocated to user $u$ to its transmission rate per PRB. In practice, $r_{u}$ is a staircase increasing function that can be well approximated as a continuous concave function, e.g., $r_{u}(P_{u})=w\log(1+P_{u}/N_{u})$ where $N_{u}$ is the noise plus interference PSD and $w$ is an appropriate constant for the data channel codes used.
We allocate power across UEs by maximizing a fairness function of the UE throughput under the EIRP constraint. To this aim, we introduce the fairness function [12]:
$f_{\alpha}(x):=\left\{\begin{array}[]{ll}\frac{x^{1-\alpha}}{1-\alpha}\quad\mathrm{if\ }\alpha\geq 0,\ \alpha\neq 1\\ \log(x)\quad\mathrm{if\ }\alpha=1\end{array}\right..$ (10)
and we formulate our power allocation problem for each segment $s\in\mathcal{S}$ and slot $k$ as follows:
$\max_{\{P_{u}\}}$ $\ \sum_{u\in\mathcal{U}}f_{\alpha}\left(A_{u}r_{u}(P_{u})\right)$ (11a)
$\mathrm{s.t.}$ $\ \sum_{u\in\mathcal{U}}A_{u}\,P_{u}\,\widehat{G}_{u}\leq b$ (11b)
$\ \underline{P}_{u}\leq P_{u}\leq\overline{P}_{u},\quad\forall\,u\in\mathcal{U}.$ (11c)
where in (11b) we used the consumption upper bound (9).
Note that, if $\alpha=1$, then the objective function (11a) maximizes the proportional fairness of the throughput across UEs; as $\alpha$ grows, it tends to max-min fairness [12]. The lower power value $\underline{P}_{u,k}$ can be chosen such that reception is possible with the most robust modulation and coding scheme available.
Water-filling solution. Problem (III-B) is convex, with separable objective function and subject to linear constraints. Hence, if the feasibility region contains an interior point (known as Slater condition [13]), which holds if $\underline{P}_{u}$’s are sufficiently low, then Karush-Kuhn-Tucker (KKT) conditions are necessary and sufficient conditions for optimality [13]. In our case, KKT conditions lead to the following water-filling type solution.
Define the class of power allocations $\{P_{u}^{*}(\nu)\}_{u}$, depending on the value of a parameter (read, Lagrangian multiplier) $\nu$:
$P_{u}^{*}(\nu)\!=\!\left\{\begin{array}[]{ll}\!\!\!\!\overline{P}_{u},\quad\mathrm{if}\ (A_{u}\widehat{G}_{u})^{-1}\frac{d}{dP_{u}}f_{\alpha}(A_{u}r_{u}(\overline{P}))>\nu\\ \!\!\!\!\underline{P},\quad\mathrm{if}\ (A_{u}\widehat{G}_{u})^{-1}\frac{d}{dP_{u}}f_{\alpha}(A_{u}r_{u}(\underline{P}_{u}))<\nu\\ \!\!\!\!P^{*}:\ (A_{u}\widehat{G}_{u})^{-1}\frac{d}{dP_{u}}f_{\alpha}(A_{u}r_{u}(P^{*}))=\nu,\ \mathrm{else}\end{array}\right.$ (12)
Then, the solution to the problem (III-B) is $P_{u}^{*}=P_{u}^{*}(\nu^{*})$ for all UEs $u$, where $\nu^{*}$ is the “water level” such that:
$\nu^{*}=\max\left\{\nu:\ \sum_{u}A_{u}\widehat{G}_{u}P_{u}^{*}(\nu)\leq b\right\}.$ (13)
Note that computing $\nu^{*}$ in (13) requires one to find the root of the function $\nu\rightarrow\sum_{u}A_{u}\widehat{G}_{u}P_{u}^{*}(\nu)-b$, which can be solved numerically via, e.g., the classic bisection method.
IV Simulations
Our simulation experiments are performed in a DL system-level simulator implementing a 3GPP calibrated Urban Macro (UMa) [14] channel model, abstracting the physical-layer effects through a link to system-level interface computing equivalent Signal to Interference plus Noise Ratio (SINR) at transmission time, given the cell/user topology and the beamformed active transmissions in each cell. The simulation environment consists of a hexagonal deployment of seven three-sector sites at $500$m inter-site distance, corresponding to 21 cells with gNB installed at 25m height and 12 degrees mechanical downtilt. The main simulation parameters can be found in Table I. The central frequency is at 3.5 GHz, with 273 PRBs at 30 kHz subcarrier spacing generating a 100 MHz carrier in Time Division Duplex (TDD) split, with an average ratio of 4 DL slots every uplink (UL) slot and a total simulation time of 12 seconds. In our considered massive MIMO scenario the gNBs is equipped with a uniform rectangular array of 192 cross-polarized antennas with 5.2 dBi gain. The UEs have 2 cross-polarized omni-directional antenna elements. The antenna spacing is defined in multiple of the wavelength $\lambda$ in Table I. The wideband Channel Quality Indicator (CQI) report is provided every 160 ms to the gNB, with measurements taken every 80 ms. DL Beamforming is based on Sounding Reference Signal (SRS) transmitted by the UEs in UL, with beams selected among a codebook of Discrete Fourier Transform (DFT) beams [15] in the two-dimensional angular $(\phi,\theta)$ plane. The UEs generate traffic according to the 3GPP File Transfer Protocol 2 (FTP2) traffic model [16], where each UE starts downloading a new packet of fixed dimension $Q$ exactly 50 ms after its previous packet was successfully downloaded. The MAC scheduler architecture is depicted in Fig. 1. At most 8 UEs per slot according to proportional fair metrics in the “UE selection” block, and after the possible actions due to EIRP control, finally the available PRBs are allocated to them with a round robin criterion with single-user MIMO allocations for up to 4 layers. Matching the proportional-fair criterion, we chose $\alpha=1$ in (III-B) to maximize the geometric mean of throughput for the users connected the cell.
General Environment 3GPP UMa [14], no buildings
Cells deployment $7\times 3$ sector sites at $500$m distance
Traffic model 210 FTP2 [16], Reading time $50$ ms
CQI feedback Wideband reports every 160 ms
CQI/MCS Table CQI/MCS Table 2, 256 QAM [17]
Link Performance 3GPP data channel codes [17]
Central Frequency 3.5 GHz, TDD
Subcarrier Spacing 30 kHz (0.5 ms slots)
Number of PRBs, band 273, 100 MHz
Max Transmit Power, $\overline{P}$ 53 dBm, $0.73$ W/PRB
gNB Antennas, spacing 12x8 cross-pol, $0.7\lambda$ rows, $0.5\lambda$ cols
UE Antennas, spacing 2 cross-pol elements $0.5\lambda$
Spatial multiplexing Up to rank 4 Single User MIMO
User Mobility 3 km/h
TABLE I: Main Simulation Parameters
A single segment is configured for the simulations to enforce a EIRP constraint over the whole sector $\gamma^{*}=\rho\overline{\gamma}$, where $\overline{\gamma}$ is the maximum EIRP that can be radiated by the gNB array by using the full 53 dBm transmit power and maximum array gain at boresight, and the power reduction factor is $0<\rho\leq 1$. The outer-loop algorithm curbing the long-term target is implemented according to [8]. The PL techniques described in Section III are compared against the “Resource Limiting (RL)” baseline used in [4, 5], where PRBs are consecutively allocated using $P_{u,k}=\overline{P}$ to UEs in the slot as long as (9) is satisfied. The slot budget with “RL” and “PL” is computed with (5). The “PL- R” implements the budget refinement as per (6), with $\rho^{*}=0.1$ and $b^{*}=0.1\gamma^{*}$. For reference, the curve without any EIRP control and power optimization [11] is plot with the “No EIRP Control” tag.
In Fig. 2 we compare the average throughput carried by each cell without EIRP limitation versus different EIRP enforcement techniques with $\rho=1/4$. With “No EIRP Control”, we can observe how the carried load grows with increasing packet size $Q$, until the system saturation limit around 430 Mbits/s (black). Once activating EIRP control, a lower saturation point is to be expected. The RL technique (blue) saturates much earlier at 240 Mbits/s, but way above $1/4$ of the system saturation limit, due to the average reduced interference. On the other hand, the PL techniques can be much closer to the system saturation limit, achieving up to 402 Mbits/s with the “PL- R” (red solid) and the hand-tuned $\epsilon=0.9$. One can appreciate the small, but persistent gain of “PL- R” versus the simple “PL” (green), or the choice of less or more initial budget allocations in the period, with $\epsilon=1.0,0.5$ (red dashed, dotted), respectively.
In Fig. 3 we plot the average throughput perceived by each UE, defined as the ratio between the bits successfully received versus the time the UE needs to receive them. Therefore, losses of UE throughput are observed with high packet size even without any EIRP limitation, due to the resource contention among UEs. In Fig. 3 we focus at just two EIRP control techniques, the best performing “PL- R” with $\epsilon=0.9$ (red) versus the RL baseline [4, 5]. Then, we compare their performance with increasing EIRP limitations with $\rho=1/2,1/4,1/8$ (cross, circle, plus markers), respectively. As expected, with lower $\rho$ the UE throughput degrades with lower packet sizes, hence offered load. However, we can notice the improvements brought by “PL- R” technique, that are more consistent when the EIRP limit is more demanding. This is due by the superposition of two effects discussed in [10, 11]. First of all, the throughput loss scales logarithmically with the transmit power reduction. Moreover, the power optimization techniques [10, 11] applied by PL have the effect of stabilizing and reducing interference, resulting in the potentially higher peak performance, as also observed in Fig. 3 by the 5-8% gains at low loads, when comparing to the “No EIRP Control” baseline. Interestingly enough, with half of the maximum EIRP at $\rho=1/2$, the UE throughput is reduced by less than $1\%$ at mid-high loads with “PL- R”.
Finally, even if not shown for reason of space, we can confirm that all proposed EIRP control techniques discussed in this work were successful in enforcing the desired limitation.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.400100100100100200200200200300300300300400400400400Packet Size $Q$ [Mbits]Cell Throughput [Mbits/s]1.21.21.21.21.31.31.31.31.41.41.41.4400400400400420420420420440440440440
Figure 2: Average throughput carried by each cell in the system with different EIRP control techniques at constant power reduction factor $\rho=1/4$ and the reference performance without any power adaptation and EIRP control.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.40020202020404040406060606080808080100100100100120120120120Packet Size $Q$ [Mbits]UE Throughput [Mbits/s]
Figure 3: Average UE throughput - defined as ratio between received bits and the time needed to receive them - without any power adaptation and EIRP control, for our best “PL- R” technique, and the baseline “RL”, with different power reduction factors $\rho$.
V Conclusion
In this work we propose a family of techniques to spread the EIRP budget in each slot and then to allocate power and resources to UEs upon MAC scheduling slot-by-slot operations. Differently from MAC scheduling EIRP control prior art, we dynamically curb the power per UE in a computationally feasible manner. This is achieved by minimizing dependencies among different “segments” where EIRP constraints are active and by deriving a proper water-filling solution to the resulting parallel problems.
Our proposal allows to drastically reduce the communications performance losses compared to the legacy approaches that dynamically adapt the resource allocation without any power adaptation. For example, in case of $25\%$ allowed EIRP, our best proposal loses only $6.5\%$ in cell throughput, in contrast to the $44.2\%$ reduction with the legacy approaches. The minimal computational overhead and the performance achieved in our extremely detailed simulation study confirm that our technique is a standout candidate to enforce EIRP constraints in slot-by-slot operations of 5G and future 6G MAC schedulers.
Acknowledgements
The authors would like to thank Tedros Abdu for the great discussions and feedback during the evolution of this study.

### v2

@@
-EMF Mitigation via 5G and 6G Scheduling
+EMF Exposure Mitigation via Scheduling
 Abstract
-High antenna directivity allows for high throughput transmission but also increases the exposure to electromagnetic field (EMF) of the end-users.
-Health regulations impose limitations on the incident power density, that generate a negative impact on network performance.
-In this work we focus at the slot-by-slot operations of a cellular Medium Access Control (MAC) scheduler to constrain the short-term EMF exposure upon real-time resource allocation, minimizing the impacts on network performance.
-We assume that the long-term EMF exposure is controlled by a proper outer-loop technique, that is not the object of this paper.
-Due to the minimal computational complexity allowed in MAC scheduling, existing solutions allowing practical implementation are few and focused at sub-optimal approaches curbing radio resource allocation.
-Our contribution is the derivation of a computationally efficient water-filling solution to allocate power and - then - resources, with a feasible integration of the necessary algorithms in the operations of a 5G MAC scheduler.
-We finally evaluate our proposal versus the prior art approaches with system level simulations with realistic modeling of physical and MAC level cellular procedures.
-We conclude that our proposal can control EMF with considerable less impact on network performance, making it a standout candidate for 5G and future 6G MAC scheduler implementations.
+International standards bodies define Electromagnetic field (EMF) emission requirements that can be translated into control of the base station actual Effective Isotropic Radiated Power (EIRP), i.e., averaged over a sliding time window.
+In this work we show how to comply with such requirements by designing a water-filling power allocation method operating at the Medium Access Control (MAC) scheduler level.
+Our method ensures throughput fairness across users while constraining the EIRP to a value that is produced by an outer-loop procedure which is not the focus of our paper.
+The low computational complexity of our technique is appealing given the tight computational requirements of the MAC scheduler.
+Our proposal is evaluated against the prior art approaches through massive-MIMO system level simulations that include realistic modeling of physical and MAC level cellular procedures.
+We conclude that our proposal effectively mitigates EMF exposure with considerably less impact on network performance, making it a standout candidate for 5G and future 6G MAC scheduler implementations.
 Index Terms:
-EMF, RF exposure, EIRP, MAC Scheduling, Power Control.
+EMF exposure, EIRP control, 5G, 6G, MAC Scheduling, Massive-MIMO.
 I Introduction
-The high beam directivity characterizing the deployment of fifth generation (5G) cellular networks has sparked concerns regarding potential health risks associated with electromagnetic field (EMF) exposure.
-However, regulators and the research community have long been committed to thoroughly understanding and addressing any health implications of EMF exposure through continuous regulation and oversight efforts.
-The guidelines by the International Commission on Non-Ionizing Radiation Protection (ICNIRP) [1] specify the EMF exposure limits, depending on the operating radio frequency and enforced on the average power density over a sliding time window ranging from 1 to 30 minutes [1, 2].
-Such limits can also be defined country-wise.
-Ensuring the effective enforcement of such limits while minimizing the impact on the network performance will become increasingly challenging in sixth generation (6G) Multiple Inputs Multiple Outputs (MIMO) deployments where hundreds or thousands antenna elements, hence even higher beam gains than in 5G, are expected [3].
-By adopting a conservative approach, one can ensure that EMF constraints are met even in the worst case scenario where users consistently find themselves in the direction of maximum beam gain.
-This translates to limiting the Effective Isotropic Radiated Power (EIRP) of the radiating antenna array.
-Arguably, imposing real-time constraints on EIRP in cellular networks is no easy task, due to the scheduler complexity and the tight computation latency constraint.
+Fifth generation (5G) cellular network deployments are characterized by the use of massive Multiple Inputs Multiple Outputs (MIMO) techniques which ensure high user throughout but also increases Radio Frequency (RF) electromagnetic field (EMF) exposure.
+This trend is expected to continue in sixth generation (6G) extreme-MIMO deployments, where hundreds or even thousands antenna elements, with higher beam gains than in 5G, are envisioned [1].
+When assessing a base station compliance with RF EMF exposure limits [2], the actual maximum approach described in [3] takes into account the actual Effective Isotropic Radiated Power (EIRP), i.e., averaged over a time window of 6 to 30 minutes.
+This window plays a significant role in the assessment of EMF exposure from base stations, due to the variability of their EMF emissions in time and in space.
+Channel modelling studies showed that the actual EIRP threshold is typically a factor 0.25 below the configured maximum EIRP [4], but lower values can be foreseen with advanced beamforming algorithms [5].
+Imposing real-time constraints on EIRP in cellular networks requires modifying the Medium Access Control (MAC) scheduler operations.
+This is no easy task, due to the scheduler complexity and the tight computation latency constraint.
 Existing approaches [6, 7] propose power control procedures limiting the usable bandwidth; [8] optimizes precoding for MIMO systems under the assumption of full channel knowledge.
-However, this assumption is unfeasible in reality as it would require heavy inter-working between physical and Medium Access Control (MAC) layer.
-Finally, it has been shown that imposing EIRP constraints affects network performance [7], highlighting the importance of designing algorithms that minimize the performance degradation while being implementable in MAC layer real-time procedures.
+However, this assumption is unfeasible in reality as it would require heavy inter-working between physical and MAC layer.
+Finally, imposing EIRP constraints affects network performance when reaching the actual EIRP threshold [4], highlighting the importance of designing algorithms that minimize the performance degradation while being implementable in MAC layer real-time procedures.
 Our contribution.
 We consider the framework illustrated in Fig.
-1, where an “outer-loop” mechanism operates on a slow time scale, monitoring the average EIRP over a sliding window of a few minutes, as prescribed by ICNIRP.
+1, where an “outer-loop” mechanism operates on a slow time scale, monitoring the actual EIRP over a sliding window of a few minutes, as prescribed in [3].
 Such mechanism sets the EIRP budget for the next “period” which encompasses the next few hundred of slots.
@@
 The second subroutine performs a fair allocation across the UEs while fulfilling the constraint on the per-slot EIRP budget.
-Prior art [4, 5] proposed constraining resources solely in the frequency - hence PRB- domain to limit EIRP.
+Prior art [6, 7] proposed constraining resources solely in the frequency, i.e., PRB, domain to limit EIRP.
 Instead, we propose dynamically adjusting the user MCS and power to maximize the global UE throughput fairness while adhering to an EIRP constraint.
@@
 As a side-product, overall network interference will be reduced, as demonstrated in [11, 12].
-We evaluate our proposal in a highly realistic 3rd Generation Partnership Project (3GPP) calibrated simulation setup, with a full physical and MAC implementation of a MIMO cellular scenario.
+We evaluate our proposal in a highly realistic 3rd Generation Partnership Project (3GPP) calibrated simulation setup, with a full physical and MAC implementation of a massive-MIMO cellular scenario.
 The results enable us to draw conclusions regarding the potential architecture and algorithms for deployment in a MAC scheduler of a g-Node B (gNB) in cellular networks to control EIRP with minimal impact on system performance.
@@
 Slot index $k$ has been omitted in inner-loop operations for readability.
-The main focus of our work concerns the red blocks, while the other parts are either legacy MAC scheduling operations in the inner-loop, or outer-loop operations discussed in [8]
+The main focus of our work concerns the red blocks, while the other parts are either legacy MAC scheduling operations in the inner-loop, or outer-loop operations discussed in [9].
 II EMF exposure model
-We address the downlink (DL) scheduling problem of allocating PRBs and power to users while mitigating the human exposure to EMF according to the health regulations imposed by the ICNIRP.
+We address the downlink (DL) scheduling problem of allocating PRBs and power to users while mitigating the human exposure to EMF according to [3].
 We first introduce some notation.
@@
 $\mathrm{EIRP}_{k}(\phi,\theta):=\sum_{u\in\mathcal{U}_{k}}A_{u,k}\,P_{u,k}\,G_{u,k}(\phi,\theta).$ (1)
-The power density $D$ measured at distance $R$ from the antenna in the direction $(\phi,\theta)$ is proportional to the EIRP, and equals $D=\frac{\mathrm{EIRP}(\phi,\theta)}{4\pi R^{2}}$.
-ICNIRP imposes health-based limits on $D$ [1], which can be translated into an upper bound $\overline{C}$ on the average EIRP, depending on the statistical distribution of the distance of humans from the antenna [7].
-Depending on the presence and type of human activities present in the cell (schools, hospitals, etc.), different limits may need to be imposed in different directions.
-To account for this, we partition the set of all azimuth/elevation angles, defined with respect to the radiating antenna, into a set of segments segments $\mathcal{S}$, and assign a specific EIRP limit $\overline{C}_{s}$ to each segment $s\in\mathcal{S}$.
+The power density measured at distance $R$ from the antenna in the direction $(\phi,\theta)$ is proportional to the EIRP, and equals $\frac{\mathrm{EIRP}(\phi,\theta)}{4\pi R^{2}}$ in free space conditions.
+When implementing the so-called “actual maximum approach” described in [3], the actual EIRP, i.e., time averaged over a sliding window, of the base station shall not exceed a configured actual EIRP threshold $\overline{C}$.
+Depending on the environment, different thresholds may need to be configured in different directions.
+To account for this, we partition the set of all azimuth/elevation angles, defined with respect to the radiating antenna, into a set of segments $\mathcal{S}$, and assign a specific actual EIRP threshold $\overline{C}_{s}$ to each segment $s\in\mathcal{S}$.
 We then define the EIRP consumption $c_{k,s}$ as the maximum EIRP in segment $s$:
@@
 We call period $t$ a set of $K\geq 1$ consecutive slots starting from $k_{0}^{t}$, and we define $c^{t}_{s}:=\sum_{k=k_{0}^{t}}^{k_{0}^{t}+K-1}c_{k,s}$ the sum of EIRP consumption over period $t$.
-To fulfill the ICNIRP health regulations we impose that the average EIRP consumption over every sliding window of $W$ consecutive periods does not exceed $\overline{C}_{s}$, for each segment $s$:
-$\frac{1}{W}\sum_{i=0}^{W-1}c^{t-i}_{s}\leq\overline{C}_{s},\qquad\forall\,t,\ s\in\mathcal{S}.$ (3)
+To comply with [3], we impose that the actual EIRP, i.e., the average EIRP consumption over every sliding window of $W$ consecutive periods, does not exceed $\overline{C}_{s}$, for each segment $s$:
+$\frac{1}{W}\sum_{i=0}^{W-1}c^{t-i}_{s}\leq\overline{C}_{s},\qquad\forall\,t.$ (3)
 III Power control for EMF exposure mitigation
@@
 Therefore, the available slot budget at slot $k$ is
-$b_{k,s}=\max\left(\left(1-\epsilon_{s}\frac{K-k}{K}\right)\gamma_{s}-\sum_{i=1}^{k}c_{i,k},0\right).$ (5)
+$b_{k,s}=\max\left(\left(1-\epsilon_{s}\frac{K-k}{K}\right)\gamma_{s}-\sum_{i=1}^{k-1}c_{i,s},0\right).$ (5)
 The difficulty in designing the portion $\epsilon_{s}$ resides in the fact that future user requests are unknown.
@@
 We assume that $A_{u,k}$ PRBs have been pre-allocated to each active UE $u\in\mathcal{U}_{k}$ as a function of—amongst others—the amount of bits present in the buffer for each UE, the selected MCS and the corresponding rate $m_{u},r_{u}$, respectively.
-In our preferred implementation, this is performed via the method described in [11], which downgrades the UEsMCS and reduces the power accordingly, hence increasing the required PRB allocation, but without exceeding the PRBs available in the current slot.
+In our preferred implementation, this is performed via the method described in [12], which downgrades the UEsMCS and reduces the transmit power accordingly, hence increasing the required PRB allocation, but without exceeding the PRBs available in the current slot.
 As a result, the transmit power and resulting interference in neighbouring cells are reduced.
@@
 We call $\mathcal{U}_{k,s}\subset\mathcal{U}_{k}$ the set of active users served by a beam whose main lobe is in segment $s$, i.e., $\mathcal{U}_{k,s}:=\{u\in\mathcal{U}_{k}:\,s(\beta_{u,k})=s\}$.
-Then, the EIRP consumption $c_{k,s}$ depends only on users $\mathcal{U}_{k,s}$:
-$c_{k,s}=\max_{(\phi,\theta)\in s}\sum_{u\in\mathcal{U}_{k,s}}A_{u,k}\,P_{u,k}\,G_{u,k}(\phi,\theta)$ (7)
+Then, the estimated EIRP consumption $\widehat{c}_{k,s}$ depends only on users $\mathcal{U}_{k,s}$:
+$\widehat{c}_{k,s}=\max_{(\phi,\theta)\in s}\sum_{u\in\mathcal{U}_{k,s}}A_{u,k}\,P_{u,k}\,G_{u,k}(\phi,\theta)$ (7)
 and the power allocation $P_{u,k}$ for a user in $\mathcal{U}_{k,s}$ does not impact the EIRP consumption over a different segment $s^{\prime}\neq s$.
-In this case, we can optimize the user power allocation independently for each segment.
-Next, computing $c_{k,s}$ as in (7) requires the solution of an optimization problem. To circumvent it and simplify computations, we pre-compute the maximum gain for each beam:
+In this case, we can optimize the user power allocation for the current slot $k$ independently for each segment.
+Next, computing $\widehat{c}_{k,s}$ as in (7) requires the solution of an optimization problem. To circumvent it and simplify computations, we pre-compute the maximum gain for each beam:
 $\widehat{G}_{u,k}:=\max_{(\phi,\theta)\in s}G_{u,k}(\phi,\theta)$ (8)
-and we bound $c_{k,s}$ via a simple-to-compute expression:
-$c_{k,s}\leq\,\sum_{u\in\mathcal{U}_{k,s}}A_{u,k}\,P_{u,k}\,\widehat{G}_{u,k}.$ (9)
+and we bound $\widehat{c}_{k,s}$ via a simple-to-compute expression, that must then be lower or equal than the slot budget:
+$\widehat{c}_{k,s}\leq\,\sum_{u\in\mathcal{U}_{k,s}}A_{u,k}\,P_{u,k}\,\widehat{G}_{u,k}\leq b_{k,s}.$ (9)
 Note that equality in (9) holds only if the gain is maximal in the same direction for all beams serving users in $\mathcal{U}_{k,s}$.
+Moreover, the available budget in the current slot is computed in (5)-(6) with respect to the actual, i.e., not approximate, past EIRP consumption values $c_{.,s}$ that are measured in hindsight.
 Optimization problem.
@@
 Water-filling solution.
-Problem (III-B) is convex, with separable objective function and subject to linear constraints.
+Problem (11) is convex, with separable objective function and subject to linear constraints.
 Hence, if the feasibility region contains an interior point (known as Slater condition [14]), which holds if $\underline{P}_{u}$’s are sufficiently low, then Karush-Kuhn-Tucker (KKT) conditions are necessary and sufficient conditions for optimality [14].
@@
 $P_{u}^{*}(\nu)\!=\!\left\{\begin{array}[]{ll}\!\!\!\!\overline{P}_{u},\quad\mathrm{if}\ (A_{u}\widehat{G}_{u})^{-1}\frac{d}{dP_{u}}f_{\alpha}(A_{u}r_{u}(\overline{P}))>\nu\\ \!\!\!\!\underline{P},\quad\mathrm{if}\ (A_{u}\widehat{G}_{u})^{-1}\frac{d}{dP_{u}}f_{\alpha}(A_{u}r_{u}(\underline{P}_{u}))<\nu\\ \!\!\!\!P^{*}:\ (A_{u}\widehat{G}_{u})^{-1}\frac{d}{dP_{u}}f_{\alpha}(A_{u}r_{u}(P^{*}))=\nu,\ \mathrm{else}\end{array}\right.$ (12)
-Then, the solution to the problem (III-B) is $P_{u}^{*}=P_{u}^{*}(\nu^{*})$ for all UEs $u$, where $\nu^{*}$ is the “water level” such that:
+Then, the solution to the problem (11) is $P_{u}^{*}=P_{u}^{*}(\nu^{*})$ for all UEs $u$, where $\nu^{*}$ is the “water level” such that:
 $\nu^{*}=\max\left\{\nu:\ \sum_{u}A_{u}\widehat{G}_{u}P_{u}^{*}(\nu)\leq b\right\}.$ (13)
@@
 The main simulation parameters can be found in Table I.
-The central frequency is at 3.5 GHz, with 273 PRBs at 30 kHz subcarrier spacing generating a 100 MHz carrier in Time Division Duplex (TDD) split, with an average ratio of 4 DL slots every uplink (UL) slot and a total simulation time of 12 seconds.
-In our considered massive MIMO scenario the gNBs is equipped with a uniform rectangular array of 192 cross-polarized antennas with 5.2 dBi gain.
+The central frequency is at 3.5 GHz, with 273 PRBs at 30 kHz subcarrier spacing generating a 100 MHz carrier in Time Division Duplex (TDD) split, with an average ratio of 4 DL slots every uplink (UL) slot, with a total simulation time of 12 seconds.
+In our considered scenario the gNBs are equipped with a uniform rectangular array of $12\times 8$ cross-polarized antennas with 5.2 dBi gain.
 The UEs have 2 cross-polarized omni-directional antenna elements.
@@
 The wideband Channel Quality Indicator (CQI) report is provided every 160 ms to the gNB, with measurements taken every 80 ms.
-DL Beamforming is based on Sounding Reference Signal (SRS) transmitted by the UEs in UL, with beams selected among a codebook of Discrete Fourier Transform (DFT) beams [15] in the two-dimensional angular $(\phi,\theta)$ plane.
-The UEs generate traffic according to the 3GPP File Transfer Protocol 2 (FTP2) traffic model [16], where each UE starts downloading a new packet of fixed dimension $Q$ exactly 50 ms after its previous packet was successfully downloaded.
+DL Beamforming is based on Sounding Reference Signals transmitted by the UEs in UL, with beams selected among a codebook of Discrete Fourier Transform (DFT) beams [16] in the two-dimensional angular $(\phi,\theta)$ plane.
+The UEs generate traffic according to the 3GPP File Transfer Protocol 2 (FTP2) traffic model [17], where each UE starts downloading a new packet of $Q$ bits 50 ms after its previous packet was successfully downloaded.
 The MAC scheduler architecture is depicted in Fig.
 1.
-At most 8 UEs per slot according to proportional fair metrics in the “UE selection” block, and after the possible actions due to EIRP control, finally the available PRBs are allocated to them with a round robin criterion with single-user MIMO allocations for up to 4 layers.
-Matching the proportional-fair criterion, we chose $\alpha=1$ in (III-B) to maximize the geometric mean of throughput for the users connected the cell.
+At most 8 UEs per slot are selected according to proportional-fair metrics in the “UE selection” block, and after the possible actions due to actual EIRP control, finally the available PRBs are allocated to them with a round robin criterion with single-user MIMO allocations for up to 4 layers.
+In accordance to the proportional-fair criterion, we chose $\alpha=1$ in (11).
 General Environment 3GPP UMa [15], no buildings
@@
 TABLE I: Main Simulation Parameters
-A single segment is configured for the simulations to enforce a EIRP constraint over the whole sector $\gamma^{*}=\rho\overline{\gamma}$, where $\overline{\gamma}$ is the maximum EIRP that can be radiated by the gNB array by using the full 53 dBm transmit power and maximum array gain at boresight, and the power reduction factor is $0<\rho\leq 1$.
-The outer-loop algorithm curbing the long-term target is implemented according to [8].
-The PL techniques described in Section III are compared against the “Resource Limiting (RL)” baseline used in [4, 5], where PRBs are consecutively allocated using $P_{u,k}=\overline{P}$ to UEs in the slot as long as (9) is satisfied.
-The slot budget with “RL” and “PL” is computed with (5).
-The “PL- R” implements the budget refinement as per (6), with $\rho^{*}=0.1$ and $b^{*}=0.1\gamma^{*}$.
-For reference, the curve without any EIRP control and power optimization [11] is plot with the “No EIRP Control” tag.
+A single segment is configured to enforce a EIRP constraint over the whole sector $\gamma^{*}=\rho\overline{\gamma}$, where $\overline{\gamma}$ is the maximum EIRP that can be radiated by the gNB array by using the full 53 dBm transmit power and maximum array gain, and the power reduction factor is $\rho\in(0,1]$.
+The outer-loop algorithm producing the per-period EIRP budget $\gamma$ is implemented according to [9].
+The PL techniques described in Section III are compared against the “Resource Limiting (RL)” baseline used in [6, 7], where PRBs are consecutively allocated using $P_{u,k}=\overline{P}$ to UEs each slot as long as (9) is satisfied.
+The slot budget for “RL” and “PL” is computed via (5).
+The “PL- R” option implements the budget refinement as per (6), with $\rho^{*}=0.1$ and $b^{*}=\gamma^{*}/10$.
+For reference, the curve without any EIRP control and power optimization [12] is plotted with the label “No EIRP Control”.
 In Fig.
-2 we compare the average throughput carried by each cell without EIRP limitation versus different EIRP enforcement techniques with $\rho=1/4$.
-With “No EIRP Control”, we can observe how the carried load grows with increasing packet size $Q$, until the system saturation limit around 430 Mbits/s (black).
-Once activating EIRP control, a lower saturation point is to be expected.
-The RL technique (blue) saturates much earlier at 240 Mbits/s, but way above $1/4$ of the system saturation limit, due to the average reduced interference.
-On the other hand, the PL techniques can be much closer to the system saturation limit, achieving up to 402 Mbits/s with the “PL- R” (red solid) and the hand-tuned $\epsilon=0.9$.
-One can appreciate the small, but persistent gain of “PL- R” versus the simple “PL” (green), or the choice of less or more initial budget allocations in the period, with $\epsilon=1.0,0.5$ (red dashed, dotted), respectively.
+2 we compare the average cell throughput without actual EIRP control against different EMF-compliant techniques with $\rho=1/4$ (or equivalently, $-6$ dB).
+The “No EIRP Control” curve shows how the carried load grows with increasing packet size, but saturating after a certain point, reaching 439 Mbits/s (black) at packet size $Q=1.2$ Mbits.
+Once EIRP control has been activated, a lower carrier load is to be expected.
+The RL technique (blue) saturates much earlier with a carried load of 237 Mbits/s at $Q=1.2$ Mbits, but still way above $1/4$ of “No EIRP Control” performance, thanks to the reduction in interference.
+Conversely, the PL techniques can be much closer to “No EIRP Control”, achieving up to 407 Mbits/s with the “PL- R” (red solid) and the hand-tuned $\epsilon=0.9$.
+One can appreciate the slight yet consistent advantage of “PL- R” over the simple “PL” (green), along with the effects of different initial budget allocations in the period, represented by $\epsilon=1.0,0.5$ (red dashed, dotted), respectively.
 In Fig.
-3 we plot the average throughput perceived by each UE, defined as the ratio between the bits successfully received versus the time the UE needs to receive them.
-Therefore, losses of UE throughput are observed with high packet size even without any EIRP limitation, due to the resource contention among UEs.
+3 we illustrate the average throughput perceived by each UE, defined as the ratio between the bits successfully received and the time taken for their reception.
+After an initial increase in UE throughput due to the higher carried traffic in the same time, reductions are observed with large packet size even in the absence of any actual EIRP limitation, due to the resource contention among UEs and the increased interferece.
 In Fig.
-3 we focus at just two EIRP control techniques, the best performing “PL- R” with $\epsilon=0.9$ (red) versus the RL baseline [4, 5].
-Then, we compare their performance with increasing EIRP limitations with $\rho=1/2,1/4,1/8$ (cross, circle, plus markers), respectively.
-As expected, with lower $\rho$ the UE throughput degrades with lower packet sizes, hence offered load.
-However, we can notice the improvements brought by “PL- R” technique, that are more consistent when the EIRP limit is more demanding.
-This is due by the superposition of two effects discussed in [10, 11].
-First of all, the throughput loss scales logarithmically with the transmit power reduction.
-Moreover, the power optimization techniques [10, 11] applied by PL have the effect of stabilizing and reducing interference, resulting in the potentially higher peak performance, as also observed in Fig.
-3 by the 5-8% gains at low loads, when comparing to the “No EIRP Control” baseline.
-Interestingly enough, with half of the maximum EIRP at $\rho=1/2$, the UE throughput is reduced by less than $1\%$ at mid-high loads with “PL- R”.
-Finally, even if not shown for reason of space, we can confirm that all proposed EIRP control techniques discussed in this work were successful in enforcing the desired limitation.
-000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.400100100100100200200200200300300300300400400400400Packet Size $Q$ [Mbits]Cell Throughput [Mbits/s]1.21.21.21.21.31.31.31.31.41.41.41.4400400400400420420420420440440440440
-Figure 2: Average throughput carried by each cell in the system with different EIRP control techniques at constant power reduction factor $\rho=1/4$ and the reference performance without any power adaptation and EIRP control.
-000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.40020202020404040406060606080808080100100100100120120120120Packet Size $Q$ [Mbits]UE Throughput [Mbits/s]
-Figure 3: Average UE throughput - defined as ratio between received bits and the time needed to receive them - without any power adaptation and EIRP control, for our best “PL- R” technique, and the baseline “RL”, with different power reduction factors $\rho$.
+3 we focus on just two EIRP control techniques, the best performing “PL- R” with $\epsilon=0.9$ (red) versus the RL baseline [6, 7].
+Then, we compare their performance with increasing EIRP limitations with $\rho=\{-3,-6,-9\}$ dB (cross, circle, plus markers), respectively.
+As expected, as $\rho$ decreases, the UE throughput decreases with smaller packet sizes, corresponding to reduced offered load.
+However, we can notice the improvements brought by “PL- R”, that are more consistent when the power reduction factor $\rho$ is smaller.
+This is due to the superposition of two effects discussed in [11, 12].
+First of all, the throughput loss scales logarithmically with the transmitted power reduction.
+Moreover, the power optimization techniques [11, 12] applied by PL have the effect of stabilizing and reducing interference, resulting in a potentially higher peak performance.
+This trend is also evident in Fig.
+3, where we observe gains of 5-8% at low loads compared to the ’No EIRP Control’ case.
+When $\rho=-3$ dB, the UE throughput experiences a reduction of less than $1\%$ at mid-high loads with “PL- R”.
+All proposed EIRP control techniques discussed in this work successful enforced the desired EIRP limitation, although we do not provide explicit evidence due to space constraints.
+[…]
+Figure 2: Average cell throughput with different EIRP control techniques at constant $\rho=-6$ dB, benchmarked against the achievable performance without any EIRP control.
+[…]
+Figure 3: Average UE throughput (defined as ratio between received bits and the time needed to receive them) without any power adaptation and EIRP control, for our best performing “PL- R” technique, and the baseline “RL”, as the power reduction factor $\rho$ varies.
 V Conclusion
-In this work we propose a family of techniques to spread the EIRP budget in each slot and then to allocate power and resources to UEs upon MAC scheduling slot-by-slot operations.
-Differently from MAC scheduling EIRP control prior art, we dynamically curb the power per UE in a computationally feasible manner.
-This is achieved by minimizing dependencies among different “segments” where EIRP constraints are active and by deriving a proper water-filling solution to the resulting parallel problems.
-Our proposal allows to drastically reduce the communications performance losses compared to the legacy approaches that dynamically adapt the resource allocation without any power adaptation.
-For example, in case of $25\%$ allowed EIRP, our best proposal loses only $6.5\%$ in cell throughput, in contrast to the $44.2\%$ reduction with the legacy approaches.
-The minimal computational overhead and the performance achieved in our extremely detailed simulation study confirm that our technique is a standout candidate to enforce EIRP constraints in slot-by-slot operations of 5G and future 6G MAC schedulers.
+In this work we propose a set of techniques for EMF exposure mitigation that control the base station EIRP on a slot-per-slot basis to adhere to a budget defined over multiple slots.
+This budget is produced by an outer-loop algorithm such as the one in [9].
+The integration of these two techniques enables a base station to comply with the “actual maximum approach” defined in [3].
+Our solution reduces the transmission power fairly across all UEs with relatively low computational complexity.
+Moreover, it is fully parallelizable across different segments, that partition the sector into azimuth/elevation angle clusters on which configurable EIRP thresholds.
+Our proposal significant alleviates the impact of EMF mitigation constraints on user performance compared to legacy approaches that only adjust resource allocation without any power adaptation.
+For example, when the power reduction factor is set to $1/4$ (or $-6$ dB), our method reduces cell throughput at high loads by a mere $6.5\%$, in contrast to a $44.2\%$ reduction with the legacy approaches.
+The minimal computational overhead and the performance achieved in our detailed massive-MIMO simulation study confirm that our technique is a standout candidate for enforcing actual EIRP constraints in slot-by-slot operations of 5G and future 6G MAC schedulers.
 Acknowledgements
-The authors would like to thank Tedros Abdu for the great discussions and feedback during the evolution of this study.
+The authors thank Tedros Abdu and Alois Herzog for the discussions and feedback during the evolution of this study.